Electronic Journal of Qualitative Theory of Differential Equations 2004, No. 7, 1-1; http://www.math.u-szeged.hu/ejqtde/

Addendum to Alternative Analysis Generated by a Differential Equation

Z. Z. Khukhunashvili, V. Z. Khukhunashvili Department of Mathematics and Computer Science,

Tougaloo College, Tougaloo, MS 39149 Department of Mathematics, Tbilisi State University,

Tbilisi, Georgia

February 24 2004

Abstract

This addendum concerns the paper of the above title found in EJQTDE No. 2 (2003 ). On page 3, before (2.2), instead of ”Applying the mean value theorem” it should read ”By simple calculations”. It does not change any results presented in the paper, but excludes the mistake of applying the mean value theorem for a complex valued function of a real variable. The mean value theorem is not needed.

EJQTDE, 2004 No. 7, p. 1